The diameter of an open-topped tank 1.5 m high increases uniformly from 4.2 m at the [20] base to 6 m at the top. Discharge from the base of the tank takes place through a

75 mm diameter pipe which is 3 m in length. The pipe discharges the water to the atmosphere 1.5 m below the base of the tank. Initially, the tank is fully filled with water.

Take ƒ = 0.01. Find the time to empty the tank.


1
Expert's answer
2021-10-18T03:44:25-0400


dp=75 mm= 0.075m

lp=3m

d1=4.2m, r1=2.1m

d2=6m, r2=3m

h1=1.5m

h2="\frac{0.7*1.5}{0.3}=3.5m"

Volume of water in tank,

V="\frac{1}{3}[π{r_2}^{2}(h_1+h_2)-π{r_1}^{2}h_2]"

="\frac{1}{3}[π*3^{2}*(1.5+3.5)-π*2.1^{2}*3.5]=30.95m^{2}"

Area of pipe,

"=\frac{π}{4}d^{2}=\frac{π}{4}*0.075^{2}=4.416*10^{-3}m^{2}"

Apply Bernoulli's equation;

"\frac{p_1}{\rho g}+\frac{{V_1}^{2}}{2g}+h_1+h_3=\frac{p_2}{\rho g}+\frac{{V_2}^{2}}{2g}+0+h_f"

"h_1+h_3=\frac{{V_2}^{2}}{2g}+h_f" this is equation (i)

Where "h_f=\frac{fLV^{2}}{2gd}=\frac{0.2*3*{V_2}^{2}}{2*9.81*0.075}=0.02039{V_2}^{2}"

Substituting values to equation (i)

"1.5+1.5=\frac{{V_2}^{2}}{2*9.81}+0.02039{V_2}^{2}"

V2=6.484m/s

Area of pipe*V2="\frac{Volume\ of\ water\ in\ tank}{time}"

"4.416*10^{-3}*6.484=\frac{30.95}{t}"

"t=\frac{30.95*10^{3}}{4.416*6.484}"

t=1080.91s